OpenStudy (anonymous):

Use the properties of logarithms to expand the following logarithms completely. ln(x^3z^-7)/(y^-8)

OpenStudy (anonymous):

\[3\ln \left( x \right)-7\ln \left( z \right)+8\ln \left( y \right)\]